Calculate the molar concentration of oxygen in water at 25oC for a partial pressure of 

0.22 atm. The Henry’s law of constant for oxygen is 1.3x10-3 mol/L*atm


1
Expert's answer
2021-03-16T08:25:12-0400

According to Henry's law P = KH × x


Hence XO2 = PO2 /KH "=\frac{0.22}{1.3×10^{-3}}= 169.23"


Moles of water "= \frac{1000}{18}= 55.5mol"


"Morality = 169.23 \u00d7 55.5=9392.265M"